2. Determine the angle u between BA and BC 2 m 4 m F = 3 kN
Answer #1

BA = A - B = (0, -2 , 0) - (0, 0, 0 ) = -2 j

BC = C - B = (2, 4, -1) - (0, 0, 0) = 2i + 4j - k


using dot product,

BA.BC = |BA||BC| cos(theta)

(-2 x 4) = 2 x sqrt(2^2 + 4^2 + 1^1) cos(theta)

cos(theta) = -0.878

theta = 151 deg